I have a problem I can't figure out , please help
x-(15x-6)=104
just distribute the the negative sign through out the parenthesese. like so
x-15x+6=104
Then put the two x's together like so
-14x+6=104
Then do 104 minus 6 which equals 98 so now the problem looks like
-14x=98
then divide 98 by -14 and get an answer of
-7 which then means x=-7

Found 2 solutions by sukusu, as09cheer:
Answer by sukusu(21)

ok...
x-15x+6=104
14x+6=104
14x=98
divide 98 by 14
have fun

Answer by as09cheer(1)

just distribute the the negative sign through out the parenthesese. like so
x-15x+6=104
Then put the two x's together like so
-14x+6=104
Then do 104 minus 6 which equals 98 so now the problem looks like
-14x=98
then divide 98 by -14 and get an answer of
-7 which then means x=-7